Business and Finance. Use the simple interest formula to find the principal if the total interest earned was %150 and the rate of interest was 4% for 2 years.
Answer by stanbon(48510)

Use the simple interest formula to find the principal if the total interest earned was %150 and the rate of interest was 4% for 2 years.
--------------
A = P(1+r)^n
A = 150(1+0.04)^2
A = 150(1.04)^2
A = $162.24
====================
Cheers,
Stan H.